# Learned feature representations are biased by complexity, learning order, position, and more

### v1

@@ 1 Introduction
To do so, we train a variety of models, including MLPs, ResNets and Transformers, to compute multiple independent abstract features about an input, through separate outputs (Fig. 1). This approach provides a controlled setting in which each abstract input feature is playing an equivalent computational role—contributing to predicting exactly one output value. In this controlled setting, we manipulate various properties like feature complexity, the order in which the features are trained, and their prevalence and distribution within the inputs. We show that, although these properties do not substantially affect the computational role each feature plays in the input-output behavior of the model, they substantially bias the structure of the learned representations. Specifically:
We find that features that are simpler to compute (e.g. linearly computable from the input) tend to be more strongly represented by the model than complex (e.g. non-linear) features that play an equivalent computational role. We also find that features that are learned earlier tend to be represented more strongly than those that are learned later. Thus, learned representations are substantially path-dependent—if a model is pretrained to compute feature A, then trained to additionally compute feature B, it will have different representations than if it was trained on feature B before feature A, or if it was trained on both A and B simultaneously—even if all models end up with similar behavior on train and test sets. Other properties can likewise can bias feature representations, e.g. the prevalence of a feature in the training dataset, or the output position in which the model is trained to report it. We show these broad patterns of results across a range of models and optimizers; they seem to be rather general patterns of gradient-based deep representation learning rather than artifacts of a particular architecture. However, there are complex interactions between model architecture, optimizer, and task, which can shift or even reverse the biases in some cases.
We finally discuss some broader potential implications of these findings for interpreting learned representations, comparing them between different systems, or trying to alter them: in particular, we highlight the risk that conclusions or results may be biased by relatively extraneous details of the learning process. We set these observations in the broader context of related work, such as prior work on the computational- and feature-level inductive biases of deep learning, and challenges of interpretability. Our results may have implications for both artificial intelligence and cognitive neuroscience.
@@ 2 Definitions & methods
an abstract, task-relevant property of an input; e.g. whether the phrase “to be or not to be” appears within a sentence, or whether XOR of the first and third boolean value is true. Generally, the features we discuss are those a model is explicitly trained to output.
@@
the behavior (generally an input-output mapping) of a system, described at the abstract level of the task-relevant features. For example, if we train an image-processing model to report the shape and color of an object in an image, and it continues to do so even on a held-out test set of images, we say that the model is computing the features shape and color.
@@
The downstream role of a feature in the desired output behavior, once that feature’s value has been computed. For example, if the network is trained to label sentences with whether they contain the word “cake”, and whether they are grammatical, through separate outputs we would say these features play an equivalent computational role (direct output) once they are computed, even though computing the value of one feature might be much easier.
@@
the part of a representation (e.g. a vector subspace) that can be predicted from knowing the value of a particular feature for the current input. (Note that we do not here stipulate that this subspace must be causally implicated in the model’s behavior, unlike, e.g., Cao, 2022. We make this choice for practical reasons, to avoid having to verify the causal role in every case throughout training. However, we do verify the causal role of the representations in some cases in §6.1.)
Representation variance explained by a feature: the proportion of the total variance across representations produced by a set of stimuli that is accounted for solely by knowing that feature’s value for each stimulus; i.e., the $R^{2}$ of a linear regression from the feature values for each stimulus onto the representations produced for those input stimuli. This is a method for quantifying the the “strength” of a feature in the representation.
In section 6 we justify some of these choices empirically, for example showing that the representations we study are causally implicated in each feature’s input-output mapping, illustrating how the representation variance explained by a feature impacts feature visualization and other interpretability methods, and examining how feature variance in the representations affects downstream learning.
@@ 2.1 Methods
To analyze the results, we extract internal representations from the trained model for each stimulus in the validation or test set, and then inspect those representations with respect to each feature. Our main analyses focus on the variance explained for each feature (as described in §LABEL:sec:terminology). To compute this quantity, we fit linear regressions whose inputs are the value of a feature, and whose outputs are the activations of each unit in the model’s representations. We measure the variance explained by a feature as the $R^{2}$ resulting from the corresponding regression, when the regression is fit using the validation set and tested on the test set. When comparing representation variance explained over the course of training, we normalize each relative to the total variance at the end of training, in order to plot all values on a consistent scale.
@@ 3 MLPs on binary features
Figure 2: Test accuracy (left) and representation variance (right) over learning in an MLP trained to output easy and hard features. The easy feature (reading out a single input) is learned quite rapidly, but the hard feature (sum of 4 inputs mod 2) is learned more slowly; nevertheless, by the end of training the model achieves perfect accuracy for both features on a held-out test set. Despite this, the easy feature occupies substantially more of the penultimate layer representation variance than the hard feature. It is particularly interesting to note that long after the easy feature has been mastered, the representation variance attributable to it continues to grow, especially when the harder feature performance is most rapidly improving. (Representation variance is reported normalized by the total representation variance at the end of training, to show the gradual increase over time. A version without this normalization is shown in Appendix 20. Lines are averages across 10 seeds, intervals are 95%-CIs.)
@@
We begin with some simple experiments on MLPs trained to compute simple binary features from vectors of 32 binary inputs. This initial setting is motivated by the original results of Hermann and Lampinen (2020) on RSA, and provides a simple domain where the complexity of relationships between input features and their outputs is clearly defined. We first train a 4-layer MLP to simultaneously output the values of two features embedded amongst other noisy inputs: (1) an easy, linear feature (reading out the value of an input unit) (2) a harder feature (the sum mod 2 of a non-overlapping set of 4 inputs). We study the representations of these input features at the penultimate layer of the model, immediately before the output logits.
In Fig. 2 we show the basic pattern of results, which illustrates many of the phenomena that we study in the remainder of this work. In the left panel, we show how model accuracy on a held-out test set changes over training for each feature (and in Appendix B.1 we present the corresponding loss curves). The easy linear feature is learned quite rapidly, almost instantly on the scale of the plot. By contrast, the hard feature is learned more slowly, and much later in training. In the right hand panel, we plot the portion of the variance in the penultimate representations of the model explained by each feature. Of course, because the easy feature is learned early in training, it initially begins to occupy more variance in the model feature representations. As the harder feature begins to be learned, it too occupies some variance in the penultimate representations. However, the easy feature dominates the hard one; the easier feature occupies far more variance. In fact, even though the easy feature is learned and generalized extremely early in training, as the hard feature begins to be learned, there appears to be some cooperative pressure that actually inflates the total variance occupied by the easy feature far beyond its initial value (note the inflection point in the easy feature variance explained). Thus, throughout training, the penultimate representations of the model are driven mostly by the easy feature, and learning the hard feature actually increases this dominance. [^1] We provide a simple colab demonstrating this result at https://gist.github.com/lampinen-dm/b6541019ef4cf2988669ab44aa82460b
@@ 3.2 Varying features and the ways they can be represented
If learning order alone does not explain the dominance of the easy feature, what aspects of feature complexity might? To explore this, we trained models on a larger set of hard features, ranging in difficulty from simple AND of groups of inputs to more complex ones. Importantly, in analyzing these models we considered another property of nonlinear features: there are more reasonable ways to compute them. For example, XOR(X,Y) can be decomposed as AND(OR(X,Y), NOT AND(X,Y)) or as OR(AND(X, NOT Y), AND(NOT X, Y)), which yield different patterns of representations (cf. Fig. A.14 in Hermann and Lampinen, 2020). By contrast, all simple ways of encoding the linear feature are linearly equivalent. Could it be that the model is strongly representing components of the hard feature, even if it is not representing the hard feature itself?
@@
We also observe biases in representations sparsity—more complex features tend to produce somewhat sparser representations; see Appendix 22.
Finally, in Appendix B.5 we show that there are qualitatively similar, graded biases when the models are trained simultaneously on a larger number of features sampled from a set of varying difficulty.
@@ 3.4 Exploring the effects of architecture and hyperparameters
How general are our findings? In Appx. B.6 we explore the effects of MLP architecture and other hyperparameters. We summarize the results here. First, deeper MLPs tend to show larger representational gaps between easy and hard features; model width has little effect. Despite the fact that nonlinearities strongly influence the structure of learned representations (Saxe et al., 2019; Alleman et al., 2024), our general pattern of findings is robust to choice of nonlinearity (tanh or rectifier).
Optimizers and dropout: There is, however, a surprising interaction between dropout and choice of optimizer at convergence. Without dropout, all optimizers show similar effects, with large gaps between easy and hard features. However, we show in Appx. B.6.1 that dropout yields strikingly different effects with different optimizers at convergence; while SGD and AdaGrad show consistent gaps between easy and hard features, other common optimizers like Adam show qualitatively different effects with dropout, and in fact show a slight hard feature bias at moderate dropout rates. However, this is only true at convergence—earlier in training, even when the hard features have been learned and generalize quite well, the easy features remain dominant. Furthermore, we also show that transformers trained on more complex datasets (see below) show more consistent feature representation biases across dropout levels. Thus, while these results show that the biases can be shifted in nontrivial ways by hyperparameter choices, we expect that easy feature biases will persist in standard training settings, where models are generally not trained to complete convergence on their tasks.
@@ 4 Transformers on sequence tasks
Sequential training: In Appx. B.7 we show that sequential training of one feature at a time also yields interesting biases, which shift over the course of training in distinct ways for the encoder and decoder.
@@ 4.1 Dyck languages: structural and token-level properties dominate in different model components
Figure 9: Representation variance explained at the final encoder layer (left), and pre-logits decoder layer (right) for transformers trained to identify semantic and structural features of sentences from Dyck-(20,10). Again, there is a bias towards simpler features at the penultimate layer; at the encoder layer, however, the most variance is explained by depth. (Because sequence length correlates with various features, this plot shows additional variance explained when controlling for sequence length; raw results are reported in Fig. 33. Bars are averages across 32 seeds, intervals are 95%-CIs.)
@@
However, in contrast to prior works (which have mostly focused on language modeling), we train transformers to output a variety of classifications of these sentences, including structure-independent token-level “semantic” features (a binary classification of the first bracket type), analogous token-level but structure-dependent features (classifying the type of the first bracket at maximum depth), or purely structural features (the maximum depth, number of root brackets, and maximum branching factor) of the sequence. These features do not capture the full structure of the sentences, but allow for more controlled experiments and analyses than doing full language modelling. As these structural tasks are more challenging, the models do not achieve perfect generalization performance to the held-out test set in every case, which allows us to assess the impact of imperfect generalization on the results (see below).
We show the basic effect of feature type on representation variance in Fig. 9. At the penultimate layer of the decoder, substantially more variance is explained by the simplest, non-structural, feature compared to the features involving structure. At the encoder layer, however, the pattern is nearly the opposite—most variance is explained by a structural feature (the maximum nesting depth of the sequence). Though the task formulation is different, this finding is broadly in keeping with prior results about how early layers of transformers tend to strongly represent depth in these structures (Yao et al., 2021; Friedman et al., 2023a). Overall, we again observe substantial biases in the representations of the features, but the nature of those biases shifts across the components of the model. Nevertheless, even aggregating across different components of the model, there are clear biases towards some features over others.
@@ 4.2 More language-like datasets reveal an additional output-range bias
We next evaluated transformers on a set of datasets that are more like natural language: sentences sampled from a grammar that yields real English sentences with complex syntactic structures (including center embedding or sequential dependencies), and a variety of semantic properties (such as valence and noun properties). See Appendix A.2.2 for full details. We trained transformers to output various judgements about these sentences, such as identifying the sentence structure, or reporting which noun is the main subject. We analyze the representations with respect to these features as above.
In Fig. 10 we show the results. There are several interesting results to note. At the final encoder layer, the representations are again dominated by the syntactic feature (sentence structure). However, at the penultimate decoder layer, the variance is dominated by two features that involve reporting the identity of the nouns that play different roles in the sentence. Reporting more general features of these nouns, such as whether they are plural or singular, does not lead to as high of representational variance. Likewise, reporting valence of the sentence (which can be positive, neutral, or negative) carries more variance than reporting whether the sentence is in present or past tense (two values). Thus, the number of possible output values a feature can take seem to strongly affect the variance its representations carry near the output—which seems reasonable in retrospect, given that each possible value that can be output is effectively a separate output logit. These patterns are also reflected in the encoder, but to a lesser extent. These results show that biases can still persist in slightly more naturalistic settings, and can also be driven by the number of possible outputs associated with a feature.

### v2

@@ 1 Introduction
To do so, we train various models—MLPs, ResNets and Transformers—to compute multiple independent abstract features about an input, through separate outputs (Fig. 1). This approach provides a controlled setting in which each abstract input feature plays an equivalent computational role: contributing to predicting exactly one output value. In this controlled setting, we manipulate various properties like feature complexity, the order in which features are trained, and their prevalence and distribution within the inputs. We show that, although these properties do not substantially affect the computational role each feature plays in the input-output behavior of the model, they substantially bias the structure of the learned representations. Specifically:
We find that features that are simpler to compute (e.g. linearly computable from the input) tend to be more strongly represented by the model than complex (e.g. non-linear) features that play an equivalent computational role. We also find that features that are learned earlier tend to be represented more strongly than those that are learned later. Thus, learned representations are substantially path-dependent (cf. Lubana et al., 2023)—if a model is pretrained to compute feature A, then trained to additionally compute feature B, it will have different representations than if it was trained on feature B before feature A, or if it was trained on both A and B simultaneously—even if all models end up with similar behavior on train and test sets. Other properties can likewise bias feature representations, e.g. the prevalence of a feature in the training dataset, or the output position in which the model is trained to report it. We show these broad patterns of results across a range of models and optimizers; they seem to be rather general patterns of gradient-based deep representation learning rather than artifacts of a particular architecture. However, there are complex interactions between architecture, optimizer, and task, which can shift or even reverse the biases in some cases.
We finally discuss some broader potential implications of these findings for interpreting learned representations, comparing them between different systems, or trying to alter them: in particular, we highlight the risk that conclusions or results may be biased by relatively extraneous details of the learning process. We set these observations in the broader context of related work, such as prior work on the computational- and feature-level inductive biases of deep learning, and challenges of interpretability. Our results may have implications for both artificial intelligence and cognitive (neuro)science.
@@ 2 Definitions & methods
an abstract, task-relevant property of an input; e.g. whether “to be or not to be” appears within a sentence, or whether XOR of the first and third boolean value is true. Generally, the features we discuss are those a model is explicitly trained to output.
@@
the behavior (i.e. input-output mapping) of a system, described at the abstract level of the task-relevant features. For example, if we train a vision model to report the shape and color of an object in an image, and it does so even on a held-out test set of images, we say that the model is computing the features shape and color.
@@
The downstream contribution of a feature to the desired output behavior, once that feature’s value has been computed. For example, if the network is trained to label sentences with whether they contain the word “cake”, and whether they are grammatical, through separate outputs we would say these features play an equivalent computational role (direct output) once they are computed, even though computing the value of one feature might be much easier.
@@
the portion of a representation (e.g. a vector subspace) that can be predicted from knowing the value of a particular feature for the current input. (Note that we do not here stipulate that this subspace must be causally implicated in the model’s behavior, unlike, e.g., Cao, 2022. We make this choice for practical reasons, to avoid verifying causality throughout training. However, we do verify the causal role of the representations in some cases in §6.1.)
Representation variance explained by a feature: the proportion of the total variance across representations produced by a set of stimuli that is accounted for solely by knowing that feature’s value for each stimulus; i.e., the $R^{2}$ of a linear regression from the feature values for each stimulus onto the representations produced for those input stimuli. (See methods for a more formal definition.) This is a method for quantifying the “strength” of a feature in the representation.
In section 6 we justify some of these choices empirically, for example showing that the representations are causally implicated in each feature’s input-output mapping, illustrating how the representation variance explained by a feature impacts interpretability methods, and examining how feature variance in the representations affects downstream learning.
@@ 2.1 Methods
To analyze the results, we extract internal representations from the trained model for each stimulus in the validation or test set, and then inspect those representations are driven by each feature. Our main analyses focus on the variance explained for each feature. To compute this quantity, we fit linear regressions whose inputs are the value of a feature, and whose outputs are the activations of each unit in the model’s representations. We measure the variance explained by a feature as the $R^{2}$ resulting from the corresponding regression, when the regression is fit using the validation set and tested on the test set.
More formally, if $\phi(I_{i})\in\mathbb{R}^{n}$ denotes the model’s learned representation vector for an input $I_{i}$, and $f_{i}$ is the corresponding feature value, then via linear regression we find the optimal linear predictor $W^{*}_{f}$ of the model’s representations on the validation set $V$ from that feature:
$W^{*}_{f}=\min_{W}\mathop{\mathbb{E}}_{(I_{i},f_{i})\in V}\left[||Wf_{i}-\phi\left(I_{i}\right)||^{2}\right]$
We then assess the variance explained by that feature relative to the total variance in the model’s representations on the test set $T$:
$R^{2}_{f}=\mathop{\mathbb{E}}_{(I_{i},f_{i})\in T}\left[1-\frac{||W^{*}_{f}f_{i}-\phi\left(I_{i}\right)||^{2}}{||\phi\left(I_{i}\right)||^{2}}\right]$
When comparing representation variance explained over the course of training, we normalize each timepoint relative to the total variance at the end of training (i.e., in the equation above, we replace the denominator $||\phi\left(I_{i}\right)||^{2}$ with the final value $||\phi_{final}\left(I_{i}\right)||^{2}$ at all timepoints), in order to plot all values on a consistent scale.
@@ 3 MLPs on binary features
Figure 2: Test accuracy (left) and representation variance (right) over learning in an MLP trained to output easy and hard features. The easy feature (reading out a single input) is learned quite rapidly, but the hard feature (sum of 4 inputs mod 2) is learned more slowly; nevertheless, by the end of training the model achieves perfect accuracy for both features on a held-out test set. Despite this, the easy feature occupies substantially more of the penultimate layer representation variance than the hard feature. It is particularly interesting to note that long after the easy feature has been mastered, the representation variance attributable to it continues to grow, especially when the harder feature performance is most rapidly improving. (Representation variance is plotted normalized by the total representation variance at the end of training, to show the gradual increase over time. See Appendix 20 for an unnormalized version. Lines are averages across 10 seeds, intervals are 95%-CIs.)
@@
We begin with some simple experiments on MLPs trained to compute binary features from vectors of 32 binary inputs. This initial setting is motivated by the original results of Hermann and Lampinen (2020) on RSA, and provides a simple domain where the complexity of relationships between input features and their outputs is clearly defined. We first train a 4-layer MLP to simultaneously output the values of two features embedded amongst other noisy inputs: (1) an easy, linear feature (reading out the value of an input unit): (2) a harder feature (the sum mod 2 of a non-overlapping set of 4 inputs). That is, suppose the input vector $I$ is composed of $k$ boolean inputs $I=[A,W,X,Y,Z,N_{1},...N_{k-5}]\in\{0,1\}^{k}$, where the first 5 will be relevant to one of the features and the remaining $N_{1}...N_{k-5}$ are purely noise that the model must learn to ignore. Then the features are defined respectively as:
$\text{EasyFeature}(I)$ $=A$
$\text{HardFeature}(I)$ $=(W+X+Y+Z)\!\!\mod 2$
We study the representations of these two features at the penultimate layer of the model, immediately before the output logits.
In Fig. 2 we show the basic pattern of results, which illustrates many of the phenomena that we study below. In the left panel, we plot model accuracy on a held-out test set over training for each feature (see Appendix B.1 for the corresponding loss curves). The easy linear feature is learned quite rapidly, almost instantly on the scale of the plot. [^1] In fact, the easy linear feature is linearly decodable with $\geq$98test accuracy from the penultimate representations of an untrained network. By contrast, the hard feature is learned more slowly. In the right hand panel, we plot the portion of the variance in the penultimate representations of the model explained by each feature. Of course, because the easy feature is learned early in training, it initially begins to occupy more variance in the model feature representations. As the harder feature begins to be learned, it too occupies some variance in the penultimate representations. However, the easy feature dominates the hard one; it occupies far more variance. In fact, even though the easy feature is learned and generalized extremely early in training, as the hard feature begins to be learned, there appears to be some cooperative pressure that actually inflates the total variance occupied by the easy feature far beyond its initial value (note the inflection point in the easy feature variance explained). Thus, throughout training, the penultimate representations of the model are driven mostly by the easy feature, and learning the hard feature actually increases this dominance. [^2] We provide a simple colab demonstrating this result at https://gist.github.com/lampinen-dm/b6541019ef4cf2988669ab44aa82460b
@@ 3.2 Varying features and the ways they can be represented
If learning order alone does not explain the dominance of the easy feature, what aspects of feature complexity might? To explore this, we trained models with the same easy feature, but paired against a larger set of hard features, ranging in difficulty from simple AND of groups of inputs to more complex ones. In analyzing these models we considered another property of nonlinear features: there are more reasonable ways to compute them. For example, XOR(X,Y) can be decomposed as AND(OR(X,Y), NOT AND(X,Y)) or as OR(AND(X, NOT Y), AND(NOT X, Y)), which yield different patterns of representations (cf. Fig. A.14 in Hermann and Lampinen, 2020). By contrast, all simple ways of encoding the linear feature are linearly equivalent. Could it be that the model is strongly representing components of the hard feature, even if it is not representing the hard feature itself?
@@
Per-unit analyses: We find similar qualitative biases, as well as interesting individual effects, when analyzing each unit in the penultimate layer individually (Appendix B.4).
Sparsity: We also find representation sparsity biases: more complex features tend to produce sparser representations (Appendix 24).
More than two features: Finally, in Appendix B.6 we show that there are qualitatively similar, graded biases when the models are trained simultaneously on more features sampled from a set of varying difficulty.
@@ 3.4 Exploring the effects of architecture and hyperparameters
How general are our findings? In Appx. B.7 we explore the effects of MLP architecture and other hyperparameters. We summarize the results here. First, deeper MLPs tend to show larger representational gaps between easy and hard features; model width has little effect. Despite the fact that nonlinearities strongly influence the structure of learned representations (Saxe et al., 2019; Alleman et al., 2024), our general pattern of findings is robust to choice of nonlinearity (tanh or rectifier).
Optimizers and dropout: There is, however, a surprising interaction between dropout and choice of optimizer at convergence. Without dropout, all optimizers show similar effects, with large gaps between easy and hard features. However, we show in Appx. B.7.1 that dropout yields strikingly different effects with different optimizers at convergence; while SGD and AdaGrad show consistent gaps between easy and hard features, other common optimizers like Adam show qualitatively different effects with dropout, and in fact show a slight hard feature bias at moderate dropout rates. However, this is only true at convergence—earlier in training, even when the hard features have been learned and generalize quite well, the easy features remain dominant. Furthermore, we also show that transformers trained on more complex datasets (see below) show more consistent feature representation biases across dropout levels. Thus, while these results show that the biases can be shifted in nontrivial ways by hyperparameter choices, we expect that easy feature biases will persist in standard training settings, where models are generally not trained to complete convergence on their tasks.
@@ 4 Transformers on sequence tasks
Sequential training: sequential training of one feature at a time also yields interesting biases, which shift over training in distinct ways for the encoder and decoder (Appx. B.8).
@@ 4.1 Dyck languages: structural and token-level properties dominate in different model components
Figure 9: Representation variance explained at the final encoder layer (left), and pre-logits decoder layer (right) for transformers trained to identify semantic and structural features of sentences from Dyck-(20,10). Again, there is a bias towards simpler features at the penultimate layer; at the encoder layer, however, the most variance is explained by depth. (Because sequence length correlates with features, this plot controls for sequence length; raw results are reported in Fig. 35. Bars are averages across 32 seeds, intervals are 95%-CIs.)
@@
However, in contrast to prior works (which have mostly focused on language modeling), we train transformers to output a variety of classifications of these sentences, including structure-independent token-level “semantic” features (a binary classification of the first bracket type), analogous token-level but structure-dependent features (classifying the type of the first bracket at maximum depth), or purely structural features (the maximum depth, number of root brackets, and maximum branching factor) of the sequence. These features do not capture the full structure of the sentences, but allow for more controlled experiments and analyses than doing full language modelling. As these structural tasks are more challenging, the models do not achieve perfect generalization performance to the held-out test set in every case, which allows us to assess the impact of imperfect generalization on the results.
We show the basic effect of feature type on representation variance in Fig. 9. At the penultimate decoder layer, substantially more variance is explained by the simplest, non-structural, feature compared to the features involving structure. At the encoder layer, however, the pattern is nearly the opposite—most variance is explained by a structural feature (the maximum nesting depth of the sequence). Though the task formulation is different, this finding is broadly in keeping with prior results about how early layers of transformers tend to strongly represent depth in these structures (Yao et al., 2021; Friedman et al., 2023a). Overall, we again observe substantial biases in the representations of the features, but the nature of those biases shifts across the components of the model. Nevertheless, even aggregating across different components of the model, there are clear biases towards some features over others.
@@ 4.2 More language-like datasets reveal an additional output-range bias
We next evaluated transformers on a set of datasets that are more like natural language: sentences sampled from a grammar that yields real English sentences with complex syntactic structures (center embedding or sequential dependencies), and semantic properties (such as valence and noun properties). See Appendix A.2.2 for details. We trained transformers to output various judgements about these sentences, such as identifying the sentence structure, or reporting which noun is the main subject. We analyze the representations with respect to these features.
In Fig. 10 we show the results. There are several interesting patterns. At the final encoder layer, the representations are again dominated by the syntactic feature (sentence structure). However, at the penultimate decoder layer, the variance is dominated by two features that involve reporting the identity of the nouns that play different roles in the sentence. Reporting more general features of these nouns, such as whether they are plural or singular, does not lead to as high of representational variance. Likewise, reporting valence of the sentence (which can be positive, neutral, or negative) carries more variance than reporting whether the sentence is in present or past tense (two values). Thus, the number of possible output values a feature can take seem to strongly affect the variance its representations carry near the output—which seems reasonable in retrospect, given that each possible value that can be output is effectively a separate output logit. These patterns are also reflected in the encoder, but to a lesser extent. These results show that biases can still persist in slightly more naturalistic settings, and can also be driven by the number of possible outputs associated with a feature.